# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ 1.7. Organization of the paper
Section 2 is focused on proving the partition function invariance in Theorem 1.4. The braid relations in Theorem 1.6 are proved in Section 3. Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.7 and 1.8. Appendix A contains some prior results from the work [CGS24] and some auxiliary technical results. The necessary details for the zero-temperature analogues in Theorems 1.10-1.12 are proved in Appendix B.
@@ 1.8. Notation and conventions
We use $\overset{d}{=}$ to indicate that two random variables are equal in distribution. Below, we define our notation for geometric, exponential, and log-inverse-gamma random variables.
@@ 1.9. Acknowledgements
This work was completed during the Columbia Math Summer Undergraduate Research Program in Summer 2025, and we acknowledge support for this program from the Columbia University math department. We thank George Dragomir for organizing this program. During the program, E.S. and X.Z. were partially supported by Ivan Corwin’s Simons Investigator Grant #929852. In addition, X.Z. was supported by Ivan Corwin’s NSF grant DMS:2246576. E.R.E. was supported by Princeton University’s Office of Undergraduate Research OURSIP Program through the Frances Lane Summer Research Fund. We also thank Ivan Corwin and Amol Aggarwal for helpful discussions related to this project.
2. Invariance of polymer partition function
In this section, we prove Theorem 1.4. The key to our proof is a matrix encoding of polymer partition functions, analogous to that developed in [NY04]. Section 2.1 describes the necessary algebraic properties of these matrices, and then Section 2.2 makes the connection to the polymer partition function and completes the proof.
@@ 2.1. Matrix encoding
In order to prove our polymer partition function invariance, we follow the matrix encoding method introduced at Noumi and Yamada in [NY04]. The key difference is that we will work with infinite matrices.
@@
The intuition behind this approach is that matrix multiplication $H(\mathbf{x}_{1})\cdots H(\mathbf{x}_{n})$ naturally encodes the partition function across a certain diagram with vertex weights given by the arrays $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as discussed in Lemma 2.6. Meanwhile, the outputs of map $E$ are more algebraically straightforward and closely related to the matrix inverse of $H$, which we prove in the next Lemma 2.2 as an infinite-dimensional analogue of [NY04, Equation 1.5].
@@
$\Delta H(\mathbf{x})\Delta E(-\mathbf{x})=E(-\mathbf{x})\Delta H(\mathbf{x})\Delta=\text{Id}_{\mathbb{Z}},$ (2.2)
where $\Delta=\operatorname{diag}((-1)^{i-1})_{i\in\mathbb{Z}}$ and $\text{Id}_{\mathbb{Z}}=\operatorname{diag}(1)_{i\in\mathbb{Z}}$.
@@
$=\sum_{i\in\mathbb{Z}}(-1)^{i-1}E_{i,i}+\sum_{i\leq j}\bigl{(}(-1)^{j-1}e^{x_{i}+\cdots+x_{j-1}}+(-1)^{j-2}e^{x_{i}+\cdots+x_{j-1}}\bigr{)}E_{i,k}$
@@
where the third equality is obtained by shifting the index of the second summand from $j$ to j1j-1 and last equality above follows by a telescoping of terms. A similar calculation shows that $E(-\mathbf{x})\Delta H(\mathbf{x})=\Delta$. This concludes the proof, since one can easily check that $\Delta^{2}=Id$. ∎
@@
For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have that
$E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})=E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.4)
@@
By direct computation and applying Lemma A.5 in the third equality below, we see that
$\quad\,E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))$
$=\left(\sum_{i\in\mathbb{Z}}e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-T_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=\sum_{i\in\mathbb{Z}}e^{-(D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})+T_{i}(\mathbf{W}_{2},\mathbf{W}_{1}))}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-D_{i}(\mathbf{W}_{2},\mathbf{W}_{1})}+e^{-T_{i+1}(\mathbf{W}_{2},\mathbf{W}_{1})}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\sum_{i\in\mathbb{Z}}e^{-(W_{1,i}+W_{2,i})}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-W_{1,i+1}}+e^{-W_{2,i}}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\left(\sum_{i\in\mathbb{Z}}e^{-W_{2,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-W_{1,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=E(-\mathbf{W}_{2})E(-\mathbf{W}_{1}).\qed$
Corollary 2.5.
For $\mathbf{W}_{1},\mathbf{W}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$H(\mathbf{W}_{1})H(\mathbf{W}_{2})=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).$ (2.5)
@@
First, we justify that this matrix product has finite entries. Because these $H$ matrices (as well as the $E$ matrices) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices. While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in the dot product. Further, the product is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices is finite valued.
@@
$=\Delta E(-\mathbf{W}_{1})^{-1}E(-\mathbf{W}_{2})^{-1}\Delta$
$=\Delta\bigl{(}E(-\mathbf{W}_{2})E(-\mathbf{W}_{1})\bigr{)}^{-1}\Delta$
@@
$=(\Delta E(-T(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)(\Delta E(-D(\mathbf{W}_{2},\mathbf{W}_{1}))^{-1}\Delta)$
$=H(T(\mathbf{W}_{2},\mathbf{W}_{1}))H(D(\mathbf{W}_{2},\mathbf{W}_{1})).\qed$
2.2. Relation to the periodic polymer
We now show the following relation between the $H$ matrices and the polymer partition function:

### v3

@@ 1.6. Organization of the paper
Section 2 develops the algebraic structure of the periodic Pitman transform, culminating in the proof of Theorem 1.1 in Section 2.2. Section 3 is focused on proving the partition function invariance in Theorem 1.4. Section 4 proves an inhomogeneous Burke property for the periodic Pitman transform and uses it to prove Theorems 1.5 and 1.6. Appendix A.1 contains some prior results from the work [CGS26] and some auxiliary technical results. The necessary details for the zero-temperature analogues in Theorems 1.9-1.10 are proved in Appendix A.2.
@@ 1.7. Notation and conventions
We use $\overset{d}{=}$ to indicate that two random variables or stochastic processes are equal in distribution. Below, we define our notation for geometric, exponential, and log-inverse-gamma random variables.
@@ 1.8. Acknowledgements
This work was completed during the Columbia Math Summer Undergraduate Research Program in Summer 2025, and we acknowledge support for this program from the Columbia University math department. We thank George Dragomir for organizing this program. During the program, E.S. and X.Z. were partially supported by Ivan Corwin’s Simons Investigator Grant #929852. In addition, X.Z. was supported by Ivan Corwin’s NSF grant DMS:2246576. E.R.E. was supported by Princeton University’s Office of Undergraduate Research OURSIP Program through the Frances Lane Summer Research Fund and by Princeton University’s Office of Undergraduate Research Undergraduate Fund for Academic Conferences through the Hewlett Foundation Fund. This material is based upon work supported by the National Science Foundation under Grant No. 2015553. We thank Ivan Corwin and Amol Aggarwal for helpful discussions related to this project. We also thank the anonymous referee for helpful comments on an earlier version of this paper that have greatly improved the presentation of our results.
2. Algebraic Properties of the Pitman Transform
The end goal of this section is to prove Theorem 1.1, which is done in in Section 2.2. We start by developing the necessary algebraic background. Versions of the following two results are proved in [CGS26] for shifted versions of the maps $D$ and $T$. We provide self-contained proofs in this section.
For this lemma and below, we define the notation
$\mathfrak{s}(\mathbf{X})=\sum_{i\in\mathbb{Z}_{N}}X_{i},\quad\text{for }\mathbf{X}\in\mathbb{R}^{\mathbb{Z}_{N}}.$
Lemma 2.1.
Let $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then,
  1. (i)
    $\mathfrak{s}(T(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$, and $\mathfrak{s}(D(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$.
  2. (ii)
    $\mathfrak{s}(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{2})$, and $\mathfrak{s}(\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2}))=\mathfrak{s}(\mathbf{X}_{1})$.
Proof.
These follow immediately from the definitions (1.2) and (1.13), since the sums are telescoping. ∎
Lemma 2.2.
Let $\mathbf{X}_{1},\mathbf{X_{2}}\in\mathbb{R}^{\mathbb{Z}_{N}}$. Then, for $i\in\mathbb{Z}_{N}$,
  1. (i)
    $D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})+T_{i}(\mathbf{X}_{1},\mathbf{X_{2}})=X_{1,i}+X_{2,i}$,
  2. (ii)
    $e^{-D_{i}(\mathbf{X}_{1},\mathbf{X_{2}})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X_{2}})}=e^{-X_{2,i}}+e^{-X_{1,i+1}}$, and
  3. (iii)
Proof.
Item (i): From the definition (1.2), we observe that
$T_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i-1}+\log\Biggl(\frac{e^{Y_{i-1}}\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{e^{Y_{i}}\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}\Biggr)$
$=X_{2,i-1}+Y_{i-1}-Y_{i}-\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}\Biggr)$
$=X_{2,i}+X_{1,i}-D_{i}(\mathbf{X}_{1},\mathbf{X}_{2}).$
Item (ii): We compute this as follows:
$e^{-D_{i}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})}+e^{-T_{i+1}(\mathbf{\mathbf{X}_{1}},\mathbf{\mathbf{X}_{2}})}$ $=e^{-X_{1,i+1}}\left(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{(i,j]}}}}\right)+e^{-X_{2,i}}\left(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
$=e^{-X_{2,i}}\left(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{e^{Y_{i}}\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{(i,j]}}}}\right)+e^{-X_{1,i+1}}\left(\frac{e^{Y_{i}}\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
$=e^{-X_{2,i}}\left(1+\frac{1-e^{\sum_{j\in\mathbb{Z}_{N}}Y_{j}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)+e^{-X_{1,i+1}}\left(1+\frac{e^{Y_{i}}(e^{\sum_{j\in\mathbb{Z}_{N}}Y_{j}}-1)}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
$=e^{-X_{2,i}}+e^{-X_{1,i+1}}+e^{-X_{2,i}}\left(\frac{1-e^{\sum_{j\in\mathbb{Z}_{N}}Y_{j}}}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}+\frac{e^{{\sum_{j\in\mathbb{Z}_{N}}}Y_{j}}-1}{\sum_{j\in\mathbb{Z}_{N}}{e^{Y_{[i,j]}}}}\right)$
$=e^{-X_{2,i}}+e^{-X_{1,i+1}}.$
Item (iii): This follows as the zero-temperature limit of Item (i). ∎
@@ 2.1. Matrix encoding
The proofs of Theorems 1.1 and 1.4 follow by using the algebraic properties above to prove certain matrix relations. For this, we adapt the framework of Noumi and Yamada [NY04] to infinite matrices which have a periodic structure.
@@
The intuition behind this approach is that matrix multiplication $H(\mathbf{x}_{1})\cdots H(\mathbf{x}_{n})$ naturally encodes the partition function across a certain diagram with vertex weights given by the arrays $\mathbf{x}_{1},\ldots,\mathbf{x}_{n}$ as discussed later in Lemma 3.1. Meanwhile, the outputs of map $E$ are more algebraically straightforward and closely related to the matrix inverse of map $H$, which we prove in the next Lemma 2.4 as an infinite-dimensional analogue of [NY04, Equation 1.5].
@@
$\Delta H(\mathbf{x})\Delta E(-\mathbf{x})=E(-\mathbf{x})\Delta H(\mathbf{x})\Delta=\mathrm{Id},$ (2.2)
where $\Delta=\operatorname{diag}((-1)^{i-1})_{i\in\mathbb{Z}}$ and $\mathrm{Id}=\operatorname{diag}(1)_{i\in\mathbb{Z}}$.
@@
$=\sum_{i\in\mathbb{Z}}(-1)^{i-1}E_{i,i}+\sum_{j\geq i}\bigl((-1)^{j-1}e^{x_{i}+\cdots+x_{j-1}}+(-1)^{j-2}e^{x_{i}+\cdots+x_{j-1}}\bigr)E_{i,j}$
@@
where the third equality is obtained by shifting the index of the second summand from $j$ to j1j-1 and last equality above follows by a telescoping of terms. A similar calculation shows that $E(-\mathbf{x})\Delta H(\mathbf{x})=\Delta$. This concludes the proof, since one can easily check that $\Delta^{2}=\mathrm{Id}$. ∎
@@
For $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have that
$E(-\mathbf{X}_{2})E(-\mathbf{X}_{1})=E(-D(\mathbf{X}_{1},\mathbf{X}_{2}))E(-T(\mathbf{X}_{1},\mathbf{X}_{2})).$ (2.4)
@@
By direct computation and applying Lemma 2.2 in the third equality below, we see that
$\quad\,E(-D(\mathbf{X}_{1},\mathbf{X}_{2}))E(-T(\mathbf{X}_{1},\mathbf{X}_{2}))$
$=\left(\sum_{i\in\mathbb{Z}}e^{-D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-T_{i}(\mathbf{X}_{1},\mathbf{X}_{2})}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)$
$=\sum_{i\in\mathbb{Z}}e^{-(D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})+T_{i}(\mathbf{X}_{1},\mathbf{X}_{2}))}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-D_{i}(\mathbf{X}_{1},\mathbf{X}_{2})}+e^{-T_{i+1}(\mathbf{X}_{1},\mathbf{X}_{2})}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\sum_{i\in\mathbb{Z}}e^{-(X_{1,i}+X_{2,i})}E_{i,i}+\sum_{i\in\mathbb{Z}}\left(e^{-X_{1,i+1}}+e^{-X_{2,i}}\right)E_{i,i+1}+\sum_{i\in\mathbb{Z}}E_{i,i+2}$
$=\left(\sum_{i\in\mathbb{Z}}e^{-X_{2,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)\left(\sum_{i\in\mathbb{Z}}e^{-X_{1,i}}E_{i,i}+\sum_{i\in\mathbb{Z}}E_{i,i+1}\right)=E(-\mathbf{X}_{2})E(-\mathbf{X}_{1}).\qed$
The following is the key step that will allow us to prove both Theorem 1.1 and Theorem 1.4.
Proposition 2.7.
For $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{R}^{\mathbb{Z}_{N}}$, we have
$H(\mathbf{X}_{1})H(\mathbf{X}_{2})=H(T(\mathbf{X}_{1},\mathbf{X}_{2}))H(D(\mathbf{X}_{1},\mathbf{X}_{2})).$ (2.5)
@@
First, we justify that this matrix product has finite non-negative real entries. Because the outputs of the matrix map $H$ (as well as matrix map $E$) are upper triangular, each individual index in the product consists of a finite sum of products of elements in the respective $H$ matrices. While each row and column has infinitely many non-zero terms, there are only a finite number of non-zero terms in each dot product. Further, the resulting matrix is upper triangular, so by repeating this argument, we can see that an arbitrary product of $H$ (or $E$) matrices has finite non-negative real entries.
@@
$=\Delta E(-\mathbf{X}_{1})^{-1}E(-\mathbf{X}_{2})^{-1}\Delta$
$=\Delta\bigl(E(-\mathbf{X}_{2})E(-\mathbf{X}_{1})\bigr)^{-1}\Delta$
@@
$=(\Delta E(-T(\mathbf{X}_{1},\mathbf{X}_{2}))^{-1}\Delta)(\Delta E(-D(\mathbf{X}_{1},\mathbf{X}_{2}))^{-1}\Delta)$
$=H(T(\mathbf{X}_{1},\mathbf{X}_{2}))H(D(\mathbf{X}_{1},\mathbf{X}_{2})).\qed$
2.2. Proof of Theorem 1.1
As observed previously, Item (iii) is immediate from the definition. We prove Items (i) and (ii) separately.
2.2.1. Proof of Item (i):
Our goal is to show that
$\mathcal{P}(\mathcal{P}(\mathbf{X}_{1},\mathbf{X}_{2}))=(T(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})),D(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2})))=(\mathbf{X}_{1},\mathbf{X}_{2}).$ (2.6)
Firstly, with the convention $Y_{\ell}=X_{1,\ell+1}-X_{2,\ell}$ and $Y^{\prime}_{\ell}=T_{\ell+1}(\mathbf{X}_{1},\mathbf{X}_{2})-D_{\ell}(\mathbf{X}_{1},\mathbf{X}_{2})$, notice that
$T_{i}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2}))$ $=D_{i-1}(\mathbf{X}_{1},\mathbf{X}_{2})+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}\Biggr)$
$=X_{1,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-2,j]}}}\Biggr)+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}\Biggr),$
and this equals $X_{1,i}$ if and only if
$\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}=\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-2,j]}}}.$
Hence, it is sufficient to prove that the quantity $\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}$ is the same for all $i\in\mathbb{Z}_{N}$. To show this, first observe that
$e^{Y^{\prime}_{i}}$ $=e^{X_{2,i}-X_{1,i+1}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i-1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}\Biggr)$ (2.7)
$\Longrightarrow e^{Y^{\prime}_{[i,j]}}$ $=e^{-Y_{[i,j]}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i-1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}\Biggr)$
$\Longrightarrow\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}$ $=\sum_{j\in\mathbb{Z}_{N}}e^{-Y_{(i,j]}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i-1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}\Biggr)$
$\Longrightarrow\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}$ $=\Biggl(\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\frac{e^{-Y_{(i,j]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}.$
Now, we note that $\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}-\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}=e^{\mathfrak{s}(Y)}-1$ for each $j\in\mathbb{Z}_{N}$. Then, multiplying and dividing by this term, we obtain
$\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-1,j]}}}$ $=\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{e^{\mathfrak{s}(Y)}-1}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\frac{e^{-Y_{(i,j]}}\bigl(\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}-\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}\bigr)}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}$ (2.8)
$=\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{e^{\mathfrak{s}(Y)}-1}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\Biggl(\frac{e^{-Y_{(i,j]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}-\frac{e^{-Y_{(i,j]}-Y_{j+1}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j+1,\ell]}}}\Biggr)$
$=\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{e^{\mathfrak{s}(Y)}-1}\Biggr)\Biggl(\frac{1}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}-\frac{e^{-\mathfrak{s}(Y)}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}\Biggr)=e^{-\mathfrak{s}(Y)},$
where in the penultimate step, we start the sum at $j=i$, note that the sum is telescoping, and use the fact that
$Y_{(i,j]}+Y_{j+1}=\begin{cases}Y_{(i,j+1]},&\text{if }j\not\equiv i-1\\ \mathfrak{s}(Y),&\text{if }j\equiv i-1\end{cases}.$
Our final expression does not depend on $i$, as we wanted to prove. Similarly, we notice that
$D_{i}(T(\mathbf{X}_{1},\mathbf{X}_{2}),D(\mathbf{X}_{1},\mathbf{X}_{2}))$ $=T_{i+1}(\mathbf{X}_{1},\mathbf{X}_{2})+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i-1,j]}}}\Biggr)$ (2.9)
$=X_{2,i}+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}\Biggr)+\log\Biggl(\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i-1,j]}}}\Biggr),$
and this equals $X_{2,i}$ if and only if
$\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}=\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{[i-1,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{(i-2,j]}}}.$
Hence, it is sufficient to prove that the quantity $\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}$ is the same for all $i\in\mathbb{Z}_{N}$. From our previous observations, we know that
$e^{Y^{\prime}_{(i,j]}}$ $=e^{-Y_{(i,j]}}\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[i+1,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}}\Biggr)\Biggl(\frac{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}}\Biggr)$ (2.10)
$\Longrightarrow\frac{\sum_{j\in\mathbb{Z}_{N}}e^{Y^{\prime}_{(i,j]}}}{\sum_{j\in\mathbb{Z}_{N}}e^{Y_{[i+1,j]}}}$ $=\Biggl(\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(i,\ell]}}\Biggr)\sum_{j\in\mathbb{Z}_{N}}\frac{e^{-Y_{(i,j]}}}{\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{[j+1,\ell]}}\sum_{\ell\in\mathbb{Z}_{N}}e^{Y_{(j,\ell]}}},$
which we have already shown to be equal to $e^{-\mathfrak{s}(Y)}$, which does not depend on $i$. This completes the proof.
2.2.2. Proof of Item (ii)
Notice that the relation $\mathcal{P}_{k}\mathcal{P}_{k+1}\mathcal{P}_{k}=\mathcal{P}_{k+1}\mathcal{P}_{k}\mathcal{P}_{k+1}$ translates into the following three equalities:
$T\Bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\biggr)=T\Bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\Bigr),$ (2.11)
$D\Bigl(T(\mathbf{X}_{k},\mathbf{X}_{k+1}),T\bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\bigr)\Bigr)=T\Bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\Bigr),$ (2.12)
$D\Bigl(D(\mathbf{X}_{k},\mathbf{X}_{k+1}),\mathbf{X}_{k+2}\Bigr)=D\Bigl(D\bigl(\mathbf{X}_{k},T(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\bigr),D(\mathbf{X}_{k+1},\mathbf{X}_{k+2})\Bigr).$ (2.13)
The last of these, equation (2.13) was proved previously as [CGS26, Proposition 5.13], recorded in Lemma A.2 of the appendix. We now use the matrix relation of Proposition 2.7 to demonstrate that (2.13) implies both (2.11) and (2.12). The main tool will be the following lemma: